ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 129, 2002

Aljadeff, E. and C. Kassel Explicit norm one elements for ring actions of finite abelian groups

99

Alon, G. and E. de Shalit On the cohomology of Drinfel’d’s p-adic symmetric domain

1

Babillot, M. On the mixing property for hyperbolic systems  61

De Shalit, E. See G. Alon

Defant, A., M. Maestre and P. Sevilla-Peris Cotype 2 estimates for spaces of polynomials on sequence spaces

291

Ding, Y. The gradient of certain harmonic functions on manifolds of almost nonnegative Ricci curvature

241

Fang, M.-L. and Y. Xu Normal families of holomorphic functions and shared values

125

Gelander, T. and A. Żuk Dependence of Kazhdan constants on generating subsets

93

Göbel, R., A. T. Paras and S. Shelah Groups isomorphic to all their non-trivial normal subgroups

21

Ishikawa, K. On finite p-groups which have only two conjugacy lengths

119

Jaikin-Zapirain, A. Finite groups of bounded rank with an almost regular automorphism

209

Joswig, M., V. Kaibel and F. Körner On the k-systems of a simple polytope

109

Kaibel, V. See M. Joswig

Kandemir, M. See O. Sh. Mukhtarov

Kassel, C. See E. Aljadeff

Kisin, M. Potential semi-stability of p-adic étale cohomology  157

Kochloukova, D. H. On the homological finiteness properties of some modules over metabelian Lie algebras

221

Konno, T. Twisted endoscopy and the generic packet conjecture  253

Körner, F. See M. Joswig

Kuruoǧlu, N. See O. Sh. Mukhtarov Laghribi, A. Certaines formes quadratiques de dimension au plus 6

et corps des fonctions en caractéristique 2

317

Leibman, A. Polynomial mappings of groups  29

Maestra, M. See A. Defant

Monnier, P. Poisson cohomology in dimension two  189

Mukhtarov, O. Sh., M. Kandemir and N. Kuruoǧlu Distribution of eigenvalues for the discontinuous boundary-value problem with functional-manypoint conditions

143

Paras, A. T. See R. Göbel

Raja, M. On dual locally uniformly rotund norms  77
Sastry, S. Boundary behaviour of BMO-qc automorphisms  373

Sevilla-Peris, P. See A. Defant

Shelah, S. See R. Göbel

Vishne, O. Generators of central simple p-algebras of degree 3  175

Xu, Y. See M.-L. Fang

Zhenhai, L. Existence for implicit differential equations with nonmonotone perturbations

363

Żuk, A. See T. Gelander